# Bridging Model Reference Adaptive Control and Data Informativity

### v1

@@ 6 Simulation Results
In this section, the proposed MRAC method is applied to an example of a highly maneuverable aircraft. We consider its unstable longitudinal dynamics [43, 44] with three states (angle of attack, pitch rate, and pitch angle) and four inputs (elevator, elevon, canard, and symmetric aileron). For better numerical conditioning, we scale the angle of attack by 100. Then, with sampling time of 0.01s, we get a system model in the form (1), with
$A_{\rm s}$ $=\begin{bmatrix}0.9810&0.9831&-0.0007\\ 0.0012&0.9737&0\\ 0&0.01&1\\ \end{bmatrix}\!,$ (50)
@@
Note that the matching equations (6) are satisfied, but there are an infinite number of solutions $(K,L)$, since $B_{\rm s}$ does not have full column rank.
During the implementation, the informative time $T^{*}$ is first initialized to $T^{*}=\infty$. At every time step $t$, we check whether the data $(U_{-}(t),X(t))$ satisfy (11), and update $T^{*}$ accordingly. Either $T^{*}\leq n+m$ or, if $T^{*}=\infty$ after $t=n+m$ time steps, we know from Lemma 5.7 that there are no gains $(K,L)$ satisfying (6). The online implementation of the input in Lemma 5.7 involves two distinct cases for generating the input-state data, where an informative set of data will be generated if there exist controller gains $(K,L)$ satisfying (6). After a set of informative data are collected, $T^{*}$ is recorded. According to Theorem 5.11, the control gains $(\hat{K},\hat{L})$ in (32) will converge to a solution of the matching equations (6).
In the following simulations, we set $\gamma=1.99$ and $\sigma=10^{2}$. We consider four simulations: (i) two simulations, denoted by $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, with reference input signals drawn from the normal distribution; (ii) two simulations, denoted by $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$ with constant reference input signals. With four random initial conditions, convergence to four different controllers is observed. This is expected from the fact that the matching equations have an infinite number of solutions.
To terminate the simulation when sufficient convergence is achieved, we set, motivated by (39), the stopping criterion
with $\varepsilon=10^{-10}$. With such criterion, the simulation $\mathcal{S}_{1}$ stops at $68.39$s, providing the final control gains
@@
$\hat{L}$ $=\begin{bmatrix}-0.1516&-0.1309&-0.7200&-0.3301\\ 0.9708&0.6646&-0.4491&0.6612\\ -0.1996&-0.1175&0.6504&0.0243\\ 0.5795&0.4495&1.2712&0.8365\end{bmatrix}\!.$
@@
Such matching error was below $10^{-3}$ at $40.05$s and below $10^{-4}$ at $55.24$s. The simulation $\mathcal{S}_{2}$ stops at $68.64$s, providing the final control gains
@@
with matching error $1.6612\times 10^{-5}$. Such error was below $10^{-3}$ at $41.38$s and below $10^{-4}$ at $56.66$s. We also verify that the informative time is $T^{*}=5$ in both $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, which implies that $5$ data samples are sufficient to achieve informativity for model reference control. Figure 1 shows the state tracking errors and the norms of the matching error for simulations $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$.
(a) State tracking errors in $\mathcal{S}_{1}$
(b) State tracking errors in $\mathcal{S}_{2}$
(c) Matching error in $\mathcal{S}_{1}$
(d) Matching error in $\mathcal{S}_{2}$
Figure 1: State tracking errors and norms of the matching error in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$.
[…]
@@
with matching error $1.6288\times 10^{-5}$. Such matching error was below $10^{-3}$ at $77.02$s and below $10^{-4}$ at $108.01$s. The simulation $\mathcal{S}_{4}$ stops at $135.03$s, providing the final control gains
@@
with matching error $1.6992\times 10^{-5}$. Such error was below $10^{-3}$ at $81.81$s and below $10^{-4}$ at $111.88$s. As in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, the informative time is $T^{*}=5$ in both $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$. Figures 2(e) and 2(f) illustrate the convergence of the inputs to constant values since the input signals $\hat{K}(t)x(t)+\hat{L}(t)r(t)$ are driven by constant reference input signals.
(a) State tracking errors in $\mathcal{S}_{3}$
(b) State tracking errors in $\mathcal{S}_{4}$
(c) Matching error in $\mathcal{S}_{3}$
(d) Matching error in $\mathcal{S}_{4}$
(e) Input signals in $\mathcal{S}_{3}$
(f) Input signals in $\mathcal{S}_{4}$
As compared to Fig. 1, the results in Fig. 2 exhibit slower convergence, which can be explained with the low excitation provided by a constant reference input. Nevertheless, it is worth mentioning that in all simulations, no matter if the reference input is drawn from a normal distribution or is constant, the sets of data never satisfy the full rank condition (4). Namely, the data collected at the informative time $T^{*}=5$ satisfy
@@
implying that the MRAC problem is solved in scenarios where the collected data do not allow unique identification of the system. In other words, this further confirms that the method proposed in the current paper imposes the least restrictive conditions on data collection compared to existing MRAC methods.
@@ 7 Conclusions
In this paper, the MRAC problem has been investigated from a novel perspective of data informativity. Different from previous results on data informativity that used data generated offline, in this paper we have taken an online perspective by verifying data informativity at every time step. A new MRAC method has been developed that ensures convergence of the adaptive gains to a solution of the matching equations, provided such a solution exists. As a preliminary problem, we have analyzed the relationship between data informativity for (unique) system identification and data informativity for model reference control, demonstrating that the former is only sufficient, but not necessary for informativity for model reference control. It becomes necessary only in specific scenarios. Motivated by these findings, we have devised an input function that switches between an adaptive controller and a term that increases the rank of the data matrix. Under the assumption that the matching equations have a solution, this input generates informative data for model reference control after a finite number of time steps. Moreover, in this situation, the adaptive gains of the controller are shown to converge to a solution of the matching equations. As compared to state-of-the-art MRAC approaches, the proposed method does not need knowledge of the input matrix and does not require uniqueness of the solution of the matching equations. A possibility for future research is to explore partial-information scenarios where only input and output data are available.

### v2

@@ 6 Simulation Results
In this section, the proposed MRAC method is demonstrated on a numerical system and on a practical example involving highly maneuverable aircraft dynamics. To validate the method under diverse data excitation conditions, we consider, for each system, four simulation scenarios: (i) two simulations with reference input signals drawn from the normal distribution; (ii) two simulations with constant reference input signals. Using four random initial conditions, convergence to four different controllers is observed. This is expected as the matching equations have an infinite number of solutions. For each simulation, we set $\gamma=1.99$ and $\sigma=10^{2}$. To terminate the simulation when sufficient convergence is achieved, we set, motivated by (41), the stopping criterion
[…]
6.1 Numerical system
Consider the controllable system with matrices
$A_{\rm s}$ $=\begin{bmatrix}-1.1&-0.85&-0.1&-0.62\\ -0.24&-0.65&0.77&-0.34\\ -1.57&-0.26&-0.36&-1.2\\ 0.33&-0.99&-0.43&0.56\end{bmatrix}\!$,
$B_{\rm s}$ $=\begin{bmatrix}0.35&0.52&0.01\\ 0.39&-0.14&1.45\\ 1.04&0.98&1.16\\ 0.13&0.34&-0.29\end{bmatrix}\!$,
being unknown for control design. Further consider the reference model with matrices
$A_{\rm m}$ $=\begin{bmatrix}-0.75&-0.32&0.24&-0.27\\ 0.15&0.66&-0.29&0.05\\ -0.53&1.88&-0.48&-0.16\\ 0.46&-0.94&-0.01&0.69\end{bmatrix}\!$,
$B_{\rm m}$ $=\begin{bmatrix}0.52&-0.86&-0.69\\ -0.14&1.2&0.67\\ 0.98&-0.86&-0.92\\ 0.34&-0.76&-0.55\end{bmatrix}\!$.
Two simulations $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ with reference input signals drawn from the normal distribution give the following results. The simulation $\mathcal{S}_{1}$ stops at $t=683$, providing the final control gains
$\hat{K}$ $=\begin{bmatrix}0.4429&0.3649&1.0418&0.4429\\ 0.3714&0.7567&-0.0279&0.3714\\ 0.1857&0.8784&-1.0139&0.1857\end{bmatrix}\!$,
$\hat{L}$ $=\begin{bmatrix}0.9638&-2.0056&-1.4847\\ 0.3575&-0.3296&-0.3435\\ -0.3213&1.3352&0.8282\end{bmatrix}\!$.
$||A_{\rm s}+B_{\rm s}\hat{K}-A_{\rm m},B_{\rm s}\hat{L}-B_{\rm m}||=1.0897\times 10^{-5}.$
The simulation $\mathcal{S}_{2}$ stops at $t=686$, providing the final control gains
$\hat{K}$ $=\begin{bmatrix}-0.5377&-1.2656&0.3802&-0.5377\\ 1.0252&1.8437&0.4132&1.0252\\ 0.5126&1.4219&-0.7934&0.5126\end{bmatrix}\!$,
$\hat{L}$ $=\begin{bmatrix}-0.3477&-0.0325&0.1576\\ 1.2318&-1.6450&-1.4384\\ 0.1159&0.6775&0.2808\end{bmatrix}$
with matching error $1.5017\times 10^{-5}$. For $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we obtain $X(6)$ in (54) and (55), respectively. Figure 1 shows the state tracking errors, the norms of the matching error and the input signals for $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$.
(a) State tracking errors in $\mathcal{S}_{1}$
(b) Norm of matching error in $\mathcal{S}_{1}$
(c) Input signals in $\mathcal{S}_{1}$
(d) State tracking errors in $\mathcal{S}_{2}$
(e) Norm of matching error in $\mathcal{S}_{2}$
(f) Input signals in $\mathcal{S}_{2}$
Figure 1: State tracking errors, norms of the matching error and input signals in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$.
$X(6)$ $=\begin{bmatrix}0.4515&0.5915&-0.5977&1.4086&-0.4482&-0.7842&0.9521\\ -1.1013&0.3756&-0.5818&0.4610&0.9848&0.0235&-1.6123\\ -0.1247&0.1476&0.2960&2.0724&1.3014&-0.2891&0.2497\\ 0.9257&2.2021&1.9527&2.1648&1.9960&-0.0567&0.1607\end{bmatrix}$, (54)
$X(6)$ $=\begin{bmatrix}-0.8873&0.4737&-0.2669&-0.2928&0.8949&-0.6439&-0.2982\\ -0.1354&-0.1576&0.1929&-0.5193&0.1309&0.2239&1.1206\\ -0.4604&0.2365&-0.4934&-0.4175&0.9198&-0.5817&0.7704\\ 0.3136&-0.0666&0.2349&-0.1310&0.5287&0.5892&-0.5182\end{bmatrix}$, (55)
$X(6)$ $=\begin{bmatrix}-0.8962&0.8273&0.1253&-1.7994&0.2101&1.5451&-2.6355\\ 0.2298&0.4897&0.6467&0.5691&-1.8402&1.9382&0.5407\\ 0.9684&1.4334&1.5077&-2.9396&0.1314&0.2842&0.0059\\ 1.1103&0.1631&0.2554&-1.5738&-2.2694&-0.1277&-1.5966\end{bmatrix}$, (56)
$X(6)$ $=\begin{bmatrix}0.1246&-0.3486&-0.7091&1.1875&0.4900&-1.6225&2.1005\\ -0.3700&0.4088&0.4716&-0.3069&0.6262&-0.1266&-4.5740\\ -1.1252&-0.2829&-0.3612&2.4710&-0.2162&-1.5610&-0.0264\\ -0.2358&0.0883&-0.9631&-0.9351&0.1379&-0.7048&0.4694\end{bmatrix}$, (57)
(a) State tracking errors in $\mathcal{S}_{3}$
(b) Norm of matching error in $\mathcal{S}_{3}$
(c) Input signals in $\mathcal{S}_{3}$
(d) State tracking errors in $\mathcal{S}_{4}$
(e) Norm of matching error in $\mathcal{S}_{4}$
(f) Input signals in $\mathcal{S}_{4}$
Two simulations $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$ with constant reference input $r(t)=[0.1\ 0.1\ 0.1\ 0.1]^{\top}\ \forall t\in\mathbb{Z}_{+}$ give the following results. The simulation $\mathcal{S}_{3}$ stops at $t=1260$, providing the final control gains
$\hat{K}$ $=\begin{bmatrix}0.1969&0.2704&0.2468&0.1969\\ 0.5354&0.8197&0.5021&0.5354\\ 0.2677&0.9099&-0.7489&0.2677\end{bmatrix}\!$,
$\hat{L}$ $=\begin{bmatrix}0.3203&-0.5672&-0.4438\\ 0.7864&-1.2885&-1.0375\\ -0.1068&0.8557&0.4813\end{bmatrix}$
with matching error $1.2286\times 10^{-5}$. The simulation $\mathcal{S}_{4}$ stops at $t=1265$, providing the final control gains
$\hat{K}$ $=\begin{bmatrix}0.3694&0.9106&-0.3437&0.3694\\ 0.4204&0.3929&0.8958&0.4204\\ 0.2102&0.6965&-0.5521&0.2102\end{bmatrix}\!$,
$\hat{L}$ $=\begin{bmatrix}0.1975&0.1462&-0.0257\\ 0.8683&-1.7641&-1.3162\\ -0.0658&0.6179&0.3419\end{bmatrix}$
with matching error $1.6208\times 10^{-5}$. For $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$, we obtain $X(6)$ in (56) and (57), respectively. Figure 2 shows the state tracking errors, the norms of the matching error and the input signals for $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$.
Despite the different level of data excitation that is evident from the plots of Figure 1 and Figure 2, we verify that the informative time is always $T^{*}=6$. This implies that, despite diverse data excitation conditions, $6$ data samples are always sufficient to achieve data informativity for model reference control.
6.2 Highly maneuverable aircraft example
@@
We consider the unstable longitudinal dynamics of a highly maneuverable aircraft [hartmann1979control, yuan2018robust] with three states (angle of attack, pitch rate, and pitch angle) and four inputs (elevator, elevon, canard, and symmetric aileron). For better numerical conditioning, we scale the angle of attack by 100. With sampling time of 0.01s, we get a system model in the form (1), with
$A_{\rm s}$ $=\begin{bmatrix}0.9810&0.9831&-0.0007\\ 0.0012&0.9737&0\\ 0&0.01&1\\ \end{bmatrix}\!$,
@@
The choice $B_{\rm m}=B_{\rm s}$ is based on the setting in [hartmann1979control, yuan2018robust]. Nevertheless, we stress on the fact that $B_{\rm s}$ is unknown: hence, even though we have $B_{\rm m}=B_{\rm s}$ in the simulation, we do not know that $B_{\rm m}=B_{\rm s}$ for control design.
Two simulations $\mathcal{S}_{5}$ and $\mathcal{S}_{6}$ with reference input signals drawn from the normal distribution give the following results. The simulation $\mathcal{S}_{5}$ stops at $68.39$s, providing the final control gains
@@
$\hat{L}$ $=\begin{bmatrix}-0.1516&-0.1309&-0.7200&-0.3301\\ 0.9708&0.6646&-0.4491&0.6612\\ -0.1996&-0.1175&0.6504&0.0243\\ 0.5795&0.4495&1.2712&0.8365\end{bmatrix}\!$,
with matching error
@@
Such matching error was below $10^{-3}$ at $40.05$s and below $10^{-4}$ at $55.24$s. The simulation $\mathcal{S}_{6}$ stops at $68.64$s, providing the final control gains
@@
with matching error $1.6612\times 10^{-5}$. Such error was below $10^{-3}$ at $41.38$s and below $10^{-4}$ at $56.66$s. We also verify that the informative time is $T^{*}=5$ in both $\mathcal{S}_{5}$ and $\mathcal{S}_{6}$, which implies that $5$ data samples are sufficient to achieve informativity for model reference control. In particular, in $\mathcal{S}_{5}$, we obtain the following set of state data
$X(5)=$
$\!\begin{bmatrix}0.3200&0.6376&0.8311&0.6256&0.4428&0.5105\\ 0.5871&0.4239&0.2900&0.0060&-0.0968&0.4126\\ -0.5974&-0.5915&-0.5873&-0.5844&-0.5843&-0.5853\end{bmatrix}\!$,
while in $\mathcal{S}_{6}$, we obtain
$X(5)=$
$\!\begin{bmatrix}0.7030&1.2024&0.8685&-0.2838&-0.9618&-1.0249\\ 0.6777&-0.0262&-1.2978&-0.6420&-0.8612&0.4623\\ -1.0029&-0.9961&-0.9964&-1.0093&-1.0158&-1.0244\end{bmatrix}\!$.
Figure 3 shows the state tracking errors, the norms of the matching error and the input signals for $\mathcal{S}_{5}$ and $\mathcal{S}_{6}$.
(a) State tracking errors in $\mathcal{S}_{5}$
(b) Norm of matching error in $\mathcal{S}_{5}$
(c) Input signals in $\mathcal{S}_{5}$
(d) State tracking errors in $\mathcal{S}_{6}$
(e) Norm of matching error in $\mathcal{S}_{6}$
(f) Input signals in $\mathcal{S}_{6}$
Figure 3: State tracking errors, norms of the matching error and input signals in $\mathcal{S}_{5}$ and $\mathcal{S}_{6}$.
The two simulations $\mathcal{S}_{7}$ and $\mathcal{S}_{8}$ with constant reference input $r(t)=[0.1\ 0.1\ 0.1\ 0.1]^{\top}\ \forall t\in\mathbb{Z}_{+}$ give the following results. The simulation $\mathcal{S}_{7}$ stops at $135.02$s, providing the final control gains
@@
with matching error $1.6288\times 10^{-5}$. Such matching error was below $10^{-3}$ at $77.02$s and below $10^{-4}$ at $108.01$s. The simulation $\mathcal{S}_{8}$ stops at $135.03$s, providing the final control gains
@@
with matching error $1.6992\times 10^{-5}$. Such error was below $10^{-3}$ at $81.81$s and below $10^{-4}$ at $111.88$s. As in $\mathcal{S}_{5}$ and $\mathcal{S}_{6}$, the informative time is $T^{*}=5$ in both $\mathcal{S}_{7}$ and $\mathcal{S}_{8}$. In particular, in $\mathcal{S}_{7}$, we obtain
$X(5)=$
$\!\begin{bmatrix}0.5869&-0.4179&-1.1217&-1.6710&-1.1897&-1.3823\\ -1.1421&-0.6244&-0.9753&0.5102&0.3004&-0.7382\\ 0.4350&0.4236&0.4174&0.4076&0.4127&0.4157\end{bmatrix}\!$$\qquad$
while in $\mathcal{S}_{8}$, we obtain
$X(5)=$
$\!\begin{bmatrix}2.0434&1.3784&0.7991&-0.3101&0.0711&0.8231\\ -0.7812&-0.1990&-1.5956&0.3143&0.5676&0.4209\\ 0.6150&0.6072&0.6052&0.5892&0.5924&0.5981\end{bmatrix}\!$$\qquad$
Figure 4 shows the state tracking errors, the norms of the matching error and the input signals for $\mathcal{S}_{7}$ and $\mathcal{S}_{8}$.
(a) State tracking errors in $\mathcal{S}_{7}$
(b) Norm of matching error in $\mathcal{S}_{7}$
(c) Input signals in $\mathcal{S}_{7}$
(d) State tracking errors in $\mathcal{S}_{8}$
(e) Norm of matching error in $\mathcal{S}_{8}$
(f) Input signals in $\mathcal{S}_{8}$
Figure 4: State tracking errors, norms of the matching error and input signals in $\mathcal{S}_{7}$ and $\mathcal{S}_{8}$.
It is worth mentioning that in all simulations, no matter if the reference input is drawn from a normal distribution or is constant, the sets of data never satisfy the full rank condition (4). Specifically, the data collected in $\mathcal{S}_{5}$$\mathcal{S}_{8}$ at the informative time $T^{*}=5$ satisfy
@@
implying that the MRAC problem is solved in scenarios where the collected data do not allow unique identification of the system. This further confirms that the method proposed in the current paper imposes the least restrictive conditions on data collection compared to existing MRAC methods.
@@ 7 Conclusions
In this paper, the MRAC problem has been investigated from a novel perspective of data informativity. Different from previous results on data informativity that used data generated offline, in this paper we have taken an online perspective by verifying data informativity at every time step. A new MRAC method has been developed that ensures asymptotic convergence of the adaptive gains to a solution of the matching equations, provided such a solution exists. As a preliminary problem, we have analyzed the relationship between data informativity for (unique) system identification and data informativity for model reference control, demonstrating that the former is only sufficient, but not necessary for informativity for model reference control. It becomes necessary only in specific scenarios. Motivated by these findings, we have devised an input function that switches between an adaptive controller and a term that increases the rank of the data matrix. Under the assumption that the matching equations have a solution, this input generates informative data for model reference control after a finite number of time steps. Moreover, in this situation, the adaptive gains of the controller are shown to converge to a solution of the matching equations. As compared to state-of-the-art MRAC approaches, the proposed method does not need knowledge of the input matrix and does not require uniqueness of the solution of the matching equations. A possibility for future research is to explore partial-information scenarios where only input and output data are available.